# From Evaluation to Enhancement: Large Language Models for Zero-Knowledge Proof Code Generation

### v1

@@ Abstract.
Zero-knowledge proofs (ZKPs) are increasingly deployed in domains such as privacy-preserving authentication, blockchain scalability, and secure finance. However, authoring ZK programs remains challenging: unlike mainstream programming, ZK development requires reasoning about finite field arithmetic, constraint systems, and gadgets, making it knowledge-intensive and error-prone. While large language models (LLMs) have demonstrated strong code generation capabilities in general-purpose languages, their effectiveness for ZK programming, where correctness hinges on both language mastery and gadget-level reasoning, remains unexplored. To address this gap, we propose ZK-Eval, a domain-specific evaluation pipeline that probes LLM capabilities at three levels: language knowledge, gadget competence, and end-to-end program generation. Our evaluation of four state-of-the-art LLMs reveals that models excel at surface-level syntax but struggle with gadget usage and semantic correctness, often yielding incorrect programs. Based on these insights, we introduce ZK-Coder, an agentic framework that augments LLMs with constraint sketching, guided retrieval, and interactive repair. Experiments on Circom and Noir show substantial gains, with success rates improving from 17.35% to 83.38% and from 32.21% to 90.05%, respectively. With ZK-Eval and ZK-Coder, we establish a foundation for systematically measuring and augmenting LLMs in ZK code generation to lower barriers for practitioners and advance trustworthy computation.
Zero-Knowledge Proofs, Large Language Models, Code Generation
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXjournalvolume: Xjournalnumber: Xarticle: Xjournalyear: 2025publicationmonth: 9copyright: noneccs: Software and its engineering Automatic programmingccs: Security and privacy Cryptographyccs: Software and its engineering Domain specific languagesccs: Computing methodologies Natural language processing
@@ 1. Introduction
Zero-knowledge proofs (ZKPs) are a powerful cryptographic primitive that allow one party to prove knowledge of a statement without revealing the underlying witness. In practice, this means proving that a computation or claim is correct while keeping sensitive inputs private. ZKPs have moved far beyond theory: they are now central to privacy-preserving authentication, blockchain scalability, and emerging applications across finance and security domains (Gupta, 2025; Lavin et al., 2024; team, 2024). As demand for trustworthy computation grows, the ability to author, test, and verify ZKP programs has become a first-class concern in the software engineering community (Takahashi et al., 2025; Hochrainer et al., 2024; Xu et al., 2025; Pailoor et al., 2023; Jiang et al., 2025a).
@@
However, authoring ZK programs is challenging. Unlike mainstream programming, which builds on the von Neumann model and imperative execution, ZKP programs boil down to specifying and proving mathematical constraints over finite fields. Its development requires reasoning about finite field arithmetic, constraint systems, and gadgets — a novel concept advocated by modern ZKP frameworks to encode reusable, domain-specific building blocks. As shown in Fig. 1 and Table 1, ZK programs do not describe how to execute a computation but how to prove its correctness by compiling constraints into arithmetic circuits. Developers must specify equations over inputs, outputs, and intermediate variables, carefully wiring gadgets to enforce correctness. Even small omissions or miswirings can yield satisfiable but unsound constraint systems, leading to errors that compilers rarely flag. This makes ZK development knowledge-intensive, error-prone, and inaccessible to non-experts, despite significant progress in proof systems and verification tools (Chaliasos et al., 2024; Pailoor et al., 2023; Takahashi et al., 2025).
Table 1. Comparison of Mainstream and ZK Programming.
@@
Purpose to Compute to Prove
@@
As a result, ZK programming demands significantly more time and expertise than mainstream software development, posing a steep barrier to entry for many practitioners. In contrast, recent advances in large language models (LLMs) have lowered this barrier in general-purpose languages by demonstrating strong code generation capabilities (Jiang et al., 2025b), achieving impressive results on benchmarks such as HumanEval (Chen et al., 2021) and MBPP (Research, 2021), and powering widely adopted tools like GitHub Copilot (GitHub, 2025) and Cursor (team, 2025). This contrast naturally raises the question: to what extent can modern LLMs assist in generating ZK code from natural language specifications? Answering this question requires more than surface-level benchmarking. Compared to programming in mainstream languages, success in ZK development hinges on two competencies: (i) knowledge of domain-specific languages (DSLs) and their toolchains, and (ii) mastery of gadgets as building blocks for constraint systems. Existing benchmarks focus only on mainstream languages and overlook these unique requirements.
To address this gap, we introduce ZK-Eval, a domain-specific evaluation pipeline that probes LLM capability at three levels: language knowledge, gadget-level competence, and end-to-end program code generation (left panel of Fig. 2). We anticipate that this design would provide a structured way to pinpoint where models succeed, where they fail, and why, rather than treating code generation as a monolithic task. Applying ZK-Eval to four state-of-the-art LLMs (including advanced reasoning models like GPT-o4-mini and GPT-o3) reveals a clear pattern. While models perform strongly on surface-level language knowledge, their accuracy drops significantly when reasoning about gadgets or assembling complete programs, often producing code that compiles but fails semantically. These results underscore both the promise of LLMs in lowering barriers to ZK development and the need for augmentations to bridge the gap between syntactic knowledge and reliable code generation. Motivated by these findings, we present ZK-Coder (right panel of Fig. 2), an agentic framework that augments LLMs with constraint sketching, retrieval, and interactive repair. By grounding generation in constraint reasoning and guided gadget usage, ZK-Coder achieves substantial gains in reliability, improving success rates from 17.35% to 83.38% on Circom and from 32.21% to 90.05% on Noir. In summary, our contributions are as follows:
@@ 2.1. Zero-Knowledge Proofs and Programming Frameworks
    Zero-Knowledge: If a statement is true, the verifier learns nothing beyond the validity of the statement itself.
These properties make ZKPs particularly useful in privacy-preserving applications (Gupta, 2025; team, 2024; Lavin et al., 2024). However, achieving these properties relies on the correct implementation of the underlying cryptographic primitives and protocols with respect to the high-level intent. Consequently, writing ZKP programs can be challenging due to the complexity of the underlying field-arithmetic mathematics and the need for specialized knowledge in cryptography (Pailoor et al., 2023; Wen et al., 2024b; Chaliasos et al., 2024).
Table 2. Tools for Programming Zero-Knowledge Proofs
Frontend DSL Compiler GitHub repository
Circom (Bellés-Muñoz et al., 2022) Circom Rust iden3/circom (1.5K)
Noir (Aztec Labs, 2025) Rust-like Rust noir-lang/noir (1.2K)
Leo (Chin et al., 2021) Rust-like Rust AleoHQ/leo (4.8K)
ZoKrates (Contributors, 2025c) Python-like Rust Zokrates/ZoKrates (1.9K)
Loosely speaking, developers need to translate their high-level intent into an arithmetic representation $R(x,w)=1$, where $x$ denotes public inputs, $w$ denotes the private witness, and $R$ is a collection of polynomial equations over field elements. To ease the development process, several programming frameworks have emerged, providing higher-level abstractions and DSLs for expressing ZKP computations (Sheybani et al., 2025; Ozdemir et al., 2022; Thomas, 2022). In Table 2, we review several representative frontend frameworks, including Circom, Noir, and others. These frameworks differ not only in syntax but also in how they expose the underlying constraint system, which directly impacts the difficulty of automatic code generation. Circom requires developers to think in terms of low-level circuit wiring and explicit constraints, which poses challenges for LLM-based generation since the model must produce correct and efficient circuit templates (Hochrainer et al., 2024; Chaliasos et al., 2025; Takahashi et al., 2025). Noir, by contrast, adopts a Rust-like syntax and enforces type safety, making it closer to conventional programming languages and thus more approachable for LLMs trained on general-purpose code corpora (Joel et al., 2024; Jiang et al., 2024b, 2025b). As a concurrent effort, there are also emerging frameworks that target general-purpose languages, such as SP1 (Labs, 2025), Nexus (Contributors, 2025a), and RISC Zero (Contributors, 2025b), which compiles Rust code into a restricted instruction set and generates ZK proofs over the execution trace within a virtual machine. However, these frameworks often come with significant performance overheads due to the complexity of instruction emulation within the virtual machine. Therefore, we consider them as an orthogonal direction to our work for reducing the burden of writing ZKP programs. Taken together, these tools illustrate the diverse design space of ZKP programming frontends, and highlight why aligning LLM-based code generation with specific framework characteristics is essential for usability and correctness.
In this work, we focus on Circom and Noir for two main reasons. First, they are the most widely adopted general-purpose ZK frontends in both academia and industry, with active communities and production deployments (e.g., Circom in Ethereum-based rollups, Noir in Aztec). This ensures that advances in LLM-assisted code generation for these languages can have immediate practical impact. Second, they represent two complementary design paradigms that are highly relevant for evaluating code generation: Circom exposes low-level, circuit-oriented constructs where the challenge for LLMs is to correctly translate high-level intent into constraint wiring, whereas Noir provides a Rust-like syntax with strong typing and higher-level abstractions, which better aligns with the training data of current LLMs. By studying both, we capture the spectrum from circuit-level specification to program-level ergonomics, allowing us to analyze how LLMs handle different abstraction levels in ZKP development.
@@ 2.2. LLM-based Code Generation
Large Language Models (LLMs) have emerged as powerful assistants in software development, enabling code generation from natural language prompts with tools like GitHub Copilot, Cursor, and model-driven agents quickly translating intent into code (Jiang et al., 2024b, 2025b). These systems are reshaping engineering workflows, aiding in writing, testing, and documenting code by drawing on vast amounts of publicly available repositories and documentation (GitHub, 2025).
Framework.
Modern LLM-based code generation typically follows an agentic framework (Zhang et al., 2024a; Wang et al., 2025), where the model acts as an intelligent assistant, interpreting user intent, retrieving relevant information, and producing code snippets or entire functions as needed. This process often involves multiple iterations, with the model refining its output based on user feedback and additional context. In this way, it can leverage off-the-shelf LLMs effectively without code-specific fine-tuning (Dong et al., 2025; Pan et al., 2025). However, to the best of our knowledge, there has been limited exploration of LLMs for code generation in the context of low-resource or DSLs such as those used in ZKPs.
Benchmark.
As an important part of LLM-based code generation research, benchmarks play a crucial role in understanding model performance and guiding future improvements. Currently, their accuracy and reasoning ability are usually evaluated on standard benchmarks on mainstream programming languages such as HumanEval (Chen et al., 2021) and MBPP (Research, 2021). However, evaluation tools for low-resource and DSLs remain limited, hindering comprehensive assessments of LLM performance in these contexts (Joel et al., 2024; Giagnorio et al., 2025). Code generation for ZKP programs exemplifies the challenges and opportunities in this space. Though benchmarks for low-resource languages (Cassano et al., 2022; Yang et al., 2023) and DSLs (Liu et al., 2023; Abukhalaf et al., 2023) exist, to our knowledge, no standardized benchmark suite specifically targets ZKP code generation, requiring tailored benchmarks and evaluation strategies.
@@ 3. Measuring LLM Capability for ZK Code Generation
As discussed in Sec. 1 and Sec. 2.1, ZK programming follows a proof-oriented model: developers translate high-level specifications into algebraic constraints over a finite field and produce proofs of satisfiability. This shift yields the workflow depicted in Fig. 1, from constraint specification to implementation by composing gadgets and finally proof generation. We therefore assess two capabilities essential for end-to-end success: ➀ language and toolchain knowledge and ➁ gadget-level competence.
@@
    Language and Toolchain Knowledge. Developers must understand syntax, typing rules, semantics of field operations, and the structure of witness assignments. Weakness in any of these areas routinely leads to programs that cannot be compiled, or programs that compile but encode an invalid relation.
@@
    Gadget-level Competence. Many basic operations in ZKP programs are not natively attainable in the programming frameworks. Instead, developers must rely on a library of pre-defined gadgets that encapsulate these operations with optimized implementations. Gadget-level competence refers to the ability to effectively utilize these gadgets in the context of ZKP development and, when necessary, implement new ones from scratch.
Taken together, these two capabilities form the foundation of end-to-end generation of ZK programs. Without reliable language and toolchain knowledge, LLMs are unable to produce compilable programs, while without gadget-level competence, the resulting programs cannot faithfully encode the intended relations. In practice, developers move between applying syntax and compilation rules to structure programs and assembling or extending gadgets to realize non-trivial constraints, and for LLMs, success also hinges on integrating these skills that bridge surface-level language fluency with the deeper ability to model arithmetic and logical constraints. Existing benchmarks, however, primarily target mainstream programming models, leaving this special integration of capabilities unexplored. This gap motivates our study: to systematically measure and evaluate LLM capability for ZK code generation. To study that, our evaluation pipeline introduces ZK-Eval, which explicitly measures both dimensions and assesses how they combine in end-to-end generation tasks.
@@ 3.1. ZK-Eval Benchmark
In this section, we present the curation of ZK-Eval. Instead of focusing solely on end-to-end program generation, we aim to measure the individual competencies that contribute to successful ZK code generation. Below, we outline the key components of ZK-Eval.
Figure 3. The Design of MCQ Benchmark Evaluating LLM’s Knowledge on ZK Languages.

### v2

@@ Abstract.
Zero-knowledge proofs (ZKPs) are increasingly deployed in domains such as privacy-preserving authentication, verifiable computation, and secure finance. However, authoring ZK programs remains challenging: unlike conventional software development, ZK programming manifests a fundamental paradigm shift from imperative computation to declarative verification. This process requires rigorous reasoning about finite field arithmetic and complex constraint systems (which is rare in common imperative languages), making it knowledge-intensive and error-prone. While large language models (LLMs) have demonstrated strong code generation capabilities in general-purpose languages, their effectiveness for ZK programming, where correctness hinges on both language mastery and constraint-level reasoning, remains unexplored. To address this gap, we propose ZK-Eval, a domain-specific evaluation pipeline that probes LLM capabilities on ZK programming at three levels: language knowledge, algebraic primitive competence, and end-to-end program generation. Our evaluation of four state-of-the-art LLMs reveals that while models demonstrate strong proficiency in language syntax, they struggle when implementing and composing algebraic primitives to specify correct constraint systems, frequently producing incorrect programs. Based on these insights, we introduce ZK-Coder, an agentic framework that augments LLMs with constraint sketching, guided retrieval, and interactive repair. Experiments with GPT-o3 on Circom and Noir show substantial gains, with success rates improving from 20.29% to 87.85% and from 28.38% to 97.79%, respectively. With ZK-Eval and ZK-Coder, we establish a new basis for systematically measuring and augmenting LLMs in ZK code generation to lower barriers for practitioners and advance privacy computing.
copyright: noneccs: Software and its engineering Automatic programmingccs: Security and privacy Cryptographyccs: Software and its engineering Domain specific languagesccs: Computing methodologies Natural language processing
@@ 1. Introduction
Zero-knowledge proofs (ZKPs) are a powerful cryptographic primitive that allow one party to prove knowledge of a statement without revealing the underlying witness. In practice, this means proving that a computation or claim is correct while keeping sensitive inputs private. ZKPs have moved far beyond theory: they are now central to privacy-preserving authentication, blockchain scalability, verifiable computation, and emerging applications across finance and security domains (Sun et al., 2024; Gupta, 2025; Lavin et al., 2024; team, 2024). As demand for trustworthy computation grows, the ability to author, test, and verify ZKP programs has become a first-class concern in the software engineering community (Takahashi et al., 2025; Hochrainer et al., 2024; Xu et al., 2025; Pailoor et al., 2023; Jiang et al., 2025a).
@@
However, authoring ZK programs is challenging. Unlike mainstream programming, which builds on the von Neumann model and imperative execution, ZKP programs boil down to specifying and proving mathematical constraints over finite fields. In essence, this represents a paradigm shift from computation to verification. Its development requires reasoning about finite field arithmetic and constraint systems. As shown in Fig. 1 and Table 1, ZK programs do not describe how to execute a computation but how to prove its correctness by compiling constraints into arithmetic circuits. Developers must specify equations over inputs, outputs, and intermediate variables, carefully wiring algebraic primitives to specify constraints enforcing correctness. Even small omissions or miswirings can yield unsound or incomplete constraint systems, leading to errors that compilers rarely flag. This makes ZK development knowledge-intensive, error-prone, and inaccessible to non-experts, despite significant progress in proof systems and verification tools (Chaliasos et al., 2024; Pailoor et al., 2023; Takahashi et al., 2025; Kolozyan et al., 2025; Chaliasos et al., 2025).
Table 1. Mainstream & ZK Program Comparison.
@@
Purpose to Compute to Prove/Verify
@@
As a result, ZK programming demands significantly more time and expertise than mainstream software development, posing a steep barrier to entry for many practitioners. In contrast, recent advances in large language models (LLMs) have lowered this barrier in general-purpose languages by demonstrating strong code generation capabilities (Jiang et al., 2025b), achieving impressive results on benchmarks such as HumanEval (Chen et al., 2021) and MBPP (Research, 2021), and powering widely adopted tools like GitHub Copilot (GitHub, 2025) and Cursor (team, 2025). This contrast naturally raises the question: to what extent can modern LLMs assist in generating ZK code from natural language specifications? Answering this question requires more than surface-level benchmarking. Compared to programming in mainstream languages, success in ZK development hinges on three competencies: (i) knowledge of domain-specific languages (DSLs) and their toolchains, (ii) mastery of algebraic primitives as atomic building blocks for constraint systems, and (iii) proficiency in modeling and specifying complex verification constraints. Existing benchmarks focus only on mainstream programming paradigms and overlook these requirements.
To measure this knowledge gap, this paper first introduces ZK-Eval, a domain-specific evaluation pipeline that probes LLM capability at three levels: language knowledge, primitive-level competence, and end-to-end program code generation (left panel of Fig. 2). We anticipate that this design would provide a structured way to pinpoint where models succeed, where they fail, and why.
As will be detailed in § 3, applying ZK-Eval to four SoTA LLMs (including advanced reasoning models like GPT-o4-mini and GPT-o3) reveals a clear pattern. While models perform strongly on language knowledge, their accuracy drops significantly (below 30%) when reasoning about primitives or assembling verification constraints, highlighting a critical gap between syntactic knowledge and end-to-end generation. These results underscore both the potential of LLMs to democratize ZK development and the need for targeted augmentations to bridge this capability gap. Motivated by these findings, we present ZK-Coder (right panel of Fig. 2), which combines three key components: (i) a sketch layer (ZKSL) that explicitly models constraints using algebraic primitives, (ii) retrieval-augmented generation to ground specifications in verified implementations, and (iii) interactive refinement to fix constraint misspecifications. We observe that ZK-Coder achieves substantial gains, improving code generation success rates from 20.29% to 87.85% on Circom and from 28.38% to 97.79% on Noir (using GPT-o3). Ablation studies further confirm that sketching, guided retrieval and interactive refinement are indispensable; removing either component causes a significant performance drop (over 10%). In summary, our contributions are as follows:
@@ 2.1. Zero-Knowledge Proofs and Programming Frameworks
    Zero-Knowledge: The verifier learns nothing beyond the validity of the statement itself.
These properties make ZKPs particularly useful in verifiable computation and privacy-preserving applications (Gupta, 2025; team, 2024; Lavin et al., 2024). However, ZK programming necessitates a fundamental paradigm shift from imperative computation to declarative verification. Rather than writing instructions to calculate a result, developers must model their high-level intent as an arithmetic representation $R(x,w)=1$ over a finite field, where $x$ denotes public inputs and $w$ denotes the private witness. This requirement to faithfully enforce logical intent through field-arithmetic constraints is notoriously arduous for humans; even minor errors in modeling can lead to under-constrained circuits (compromising soundness) or over-constrained systems (breaking completeness) (Pailoor et al., 2023; Wen et al., 2024a; Chaliasos et al., 2024; Kolozyan et al., 2025).
To mitigate this complexity, several programming frameworks have emerged, offering a diverse range of abstractions. We categorize these into two primary design paradigms:
  1. (1)
    Constraint-Oriented DSLs. Frameworks like Circom and Noir provide specialized syntax for expressing verification relations. Circom requires developers to think in terms of low-level circuit wiring and explicit signal management, which poses significant challenges for LLMs tasked with producing correct circuit templates (Hochrainer et al., 2024; Chaliasos et al., 2025). Conversely, Noir adopts a Rust-like syntax with strong typing, aligning more closely with the general-purpose coding corpora on which current LLMs are trained (Joel et al., 2024; Jiang et al., 2024a).
  2. (2)
    General-Purpose zkVMs. Emerging solutions like SP1 (Labs, 2025) and RISC0 (Contributors, 2025) allow developers to compile standard Rust code into a restricted instruction set, generating proofs over the execution trace within a virtual machine. While they abstract away the constraint system entirely, they often incur substantial performance overhead. We consider these an orthogonal approach to our work, as we focus on the direct generation of circuits.
In this work, we focus on Circom and Noir because they represent the most widely adopted frontends in both academia (Stronati et al., 2024; Aztec Labs, 2025; Ozdemir et al., 2022; Pailoor et al., 2023; Wen et al., 2024a) and industry (e.g., in Ethereum rollups and Aztec). Furthermore, they occupy two complementary points on the abstraction spectrum: Circom tests an LLM’s ability to perform circuit-level specification and constraint wiring, while Noir evaluates how effectively models can adapt conventional coding proficiency to the verification-oriented paradigm. By studying both, we capture the core technical challenges of ZKP development while ensuring our findings have immediate practical relevance.
@@ 2.2. LLM-based Code Generation
Large Language Models (LLMs) have emerged as powerful assistants in software development, enabling code generation from natural language prompts with tools like GitHub Copilot, Cursor, and model-driven agents quickly translating intent into code (Jiang et al., 2024a, 2025b). These systems are reshaping engineering workflows, aiding in writing, testing, and documenting code by drawing on vast amounts of publicly available repositories and documentation (GitHub, 2025).
Framework. Modern LLM-based code generation typically follows an agentic framework (Zhang et al., 2024a; Wang et al., 2025a), where the model acts as an intelligent assistant, interpreting user intent, retrieving relevant information, and producing code snippets or entire functions as needed. This process often involves multiple iterations, with the model refining its output based on user feedback or additional context. In this way, it can leverage off-the-shelf LLMs effectively without code-specific fine-tuning (Dong et al., 2025; Pan et al., 2025). However, to the best of our knowledge, there has been limited exploration of LLMs for code generation for low-resource programming languages (LRPL) or DSLs especially for ZKP DSLs.
Benchmark. Benchmarks are a crucial component of LLM-based code generation research, playing a vital role in understanding model performance and guiding improvements. Currently, their accuracy and reasoning ability are usually evaluated on standard benchmarks on mainstream programming languages such as HumanEval (Chen et al., 2021) and MBPP (Research, 2021). However, evaluation tools for LRPLs and DSLs remain limited, hindering comprehensive assessments of LLM performance in these contexts (Joel et al., 2024; Giagnorio et al., 2025). Code generation for ZKP programs exemplifies the challenges and opportunities in this space. Though benchmarks for LRPLs (Cassano et al., 2022; Yang et al., 2023) and DSLs (Liu et al., 2023; Abukhalaf et al., 2023) exist, to our knowledge, no standardized benchmark suite specifically targets ZKP code generation, requiring tailored benchmarks and evaluation strategies.
@@ 3. Measuring LLM Capability for ZK Code Generation
As discussed in § 1 and § 2.1, ZK programming follows a proof-oriented model: developers translate high-level intent into algebraic constraints over a finite field and produce proofs of satisfiability. This shift yields a unique programming workflow depicted in Fig. 1, where developers construct the verification relation by wiring together algebraic primitives (mathematical gates that enforce logical invariants over finite fields). While witness computation still resembles conventional imperative programming, the developer’s primary responsibility shifts to the rigorous specification of these verification constraints. To evaluate LLM proficiency in this critical domain, we assess two foundational capabilities essential for end-to-end success: ➀ language and toolchain knowledge and ➁ algebraic primitive competence.
@@
    Language and Toolchain Knowledge. Developers must master domain-specific syntax, strict typing rules, and the semantics of field operations to structure valid constraint specification and witness assignments. Weakness in any of these areas routinely leads to programs that cannot be compiled or produce an invalid relation that renders the proof meaningless.
@@
    Algebraic Primitive Competence. Algebraic primitives are the atomic building blocks of ZK constraint systems. Since arithmetic circuits lack native support for high-level operations, developers combine relational (e.g., comparators), logical (e.g., bitwise gates), and arithmetic (e.g., division, inversion) primitives to express logic. Competence in these atomic units is a prerequisite for constructing sound, higher-level application-specific constraint systems.
These capabilities constitute two core dimensions of end-to-end ZK program generation. Without language and toolchain knowledge, LLMs cannot produce compilable programs; without algebraic primitive competence, programs cannot faithfully encode intended relations. Since ZK programming requires integrating language-level fluency with deeper constraint modeling, existing mainstream benchmarks (Chen et al., 2021; Jain et al., 2024; Research, 2021) leave this integration unexplored. We introduce ZK-Eval to systematically measure both dimensions and their integration in end-to-end tasks.
@@ 3.1. ZK-Eval Benchmark
To realize this benchmark design, we structure ZK-Eval in three stages, each targeting one capability and their integration. We illustrate the design overview in Fig. 3.
Figure 3. The Three-Stage Evaluation Design of the ZK-Eval Benchmark.